.A person moves towards east for 3 m, then towards north for 4 m and then moves vertically up by 5 m. What is his distance now from the starting point?


1
Expert's answer
2020-08-10T19:54:03-0400


Using the Pythagorean theorem, from the right triangle "\triangle AOB" express "BO".

"BO = \sqrt{3^2 + 4^2} = 5"

Using the Pythagorean theorem again, from the right triangle "\triangle BOC" express "OC", which is the final distance from the starting point.


"OC = \sqrt{CB^2 + BO^2} = \sqrt{5^2 + 5^2} = 5\sqrt2"

Answer. "5\sqrt2".